Exhibit 5 Present Value of $1 at Compound Interest Periods Ova AWN 4% 472% 0.96154 0.956940 0.92456 0.915730 0.88900 0.876300Exhibit 7 Present Value of an Annuity of $1 at Compound Interest Periods 600 van WN- 4% 0.96154 1.88609 2.77509 3.62990 4.451Present Value of Bonds Payable; Premium Moss Co. issued $150,000 of five-year, 13% bonds, with interest payable semiannually,Moss Co. issued $150,000 of five-year, 13% bonds, with interest payable semiannually, at a market (effective) interest rate of 11%. Determine the present value of the bonds payable, using the present value tables in Exhibit 5 and Exhibit 7. Round to the nearest dollar.
Answer #1
Amount PV factor 5.5% Present value
Semi-annual interest 9750 7.53763 73492
Principal 150000 0.58543 87815
Total 161306
Present value of Bonds payable 161306 or 161307
Workings:
Semi-annual interest 9750 =150000*13%/2
PV factor interest 5.50% =11%/2
PV factor 5.5%:
Semi-annual interest 7.53763 =(1-(1.055)^-10)/0.055
Principal 0.58543 =1/1.055^10